1.   By using the knowledge of relationships among the various costs of production concepts fill the blank space of the following table

Quantity TC TFC TVC ATC AVC MC

0 125

  10 5

20 105

30 110

40 255

50 3

60 3


1
Expert's answer
2022-02-18T08:52:19-0500

i)TVC=TC-TFC"\therefore" (125-125=0,(5-125)=-120,(105-125)=-20,(110-125)=-15,(255-125)=130, (3-125)=-122,(3-125)=-122

so TVC values are; 0,-120,-20,-15,130,-122,-122

ii)ATC="\frac {TC}{OUTPUT}" "\therefore" ("\frac{125}{0}" )=0 ,"(\frac{5}{10})" =0.5 ,"\frac {105}{20}" =5.25, "\frac{110}{30}" =3.66, "\frac{255}{40}" =6.375, "\frac{3}{50}" =0.06, "\frac{3}{60}"=0.05

thus TVC values are; 0 , 0.5, 5.25, 3.66, 6.75, 0.06, 0.05

iii)AVC="\frac{TVC}{OUTPUT}" ="\frac{0}{0}" 0, "-\frac{120}{10}" =-12, "-\frac{20}{20}"=-1 ,"-\frac{15}{30}" =-0.5, "\frac{130}{40}" =3.25, "-\frac{122}{50}" =-2.44, "-\frac{122}{60}" =-2.033

values of AVC are;0, -12, -1, -0.5, 3.25, -2.44, -2.0333

iv) MC="\frac{change in TC}{change in output(Q)}" =("\frac{5-125}{20-10})" =-12, "(\frac{105-5}{30-20})" =10, "(\frac{110-105}{40-30})"=0.5, "(\frac{255-105}{50-40})" =14.5, "(\frac{3-255}{60-50})" =-12.2

the first and the last MC are =0

the MC values are; 0, -12, 10, 0.5, 14.5,-12.2, 0